### v1

# Cross-Modal Reconstruction Pretraining for Ramp Flow Prediction at Highway Interchanges

Abstract
Interchanges are pivotal junctions that facilitate vehicle transfers between highways. Accurate prediction of ramp traffic flow at highway interchanges is essential for proactive traffic management. However, the scarcity of real-time detectors on ramps creates a “real-time blind spot” challenge. While historical data can be obtained through Electronic Toll Collection (ETC) systems, real-time ramp flow cannot be directly used as model input. This data scarcity forces models to rely solely on mainline traffic data to infer complex ramp dynamics. To address this challenge, we propose a novel two-stage framework, the Spatio-Temporal Decoupled Autoencoder (STDAE), which leverages a cross-modal reconstruction pretraining task to mitigate missing data issues. In the pretraining stage, STDAE learns to reconstruct historical ramp flows exclusively from mainline traffic data, thereby compelling the model to capture the intrinsic spatio-temporal relationship between the mainline and ramp traffic flows. Its unique decoupled architecture, consisting of parallel spatial (SAE) and temporal (TAE) autoencoders, efficiently extracts heterogeneous features. In the downstream prediction stage, the learned representations are integrated with specific forecasting models such as GWNet to enhance prediction accuracy. Comprehensive experiments conducted on three real-world interchange datasets demonstrate that our proposed model (STDAEGWNET) consistently outperforms thirteen state-of-the-art baselines across multiple metrics and sampling intervals. Importantly, our method achieves performance comparable to or even surpassing advanced models that directly utilize historical ramp data, thereby validating its effectiveness in overcoming the real-time data gap. The proposed framework provides a robust and practical solution for fine-grained traffic management in detector-scarce scenarios, and its architecture-agnostic nature enables it to serve as a plug-and-play enhancement module to improve diverse forecasting pipelines. Codes and datasets are available at https://github.com/ChaochaoSeu/RPaI-STDAE.
keywords:
Traffic Flow Prediction , Highway Interchange , Spatial-Temporal Autoencoder , Pretraining , ETC Data
journal: Arxiv
[…]
[seut] organization=School of Transportation, Southeast University, city=Nanjing, postcode=211189, country=China
[…]
[zjugsc] organization=Zhejiang Urban Governance Studies Center, city=Hangzhou, postcode=311121, country=China
[…]
[nwpu] organization=School of Aeronautics, Northwestern Polytechnical University, city=Xi’an, postcode=710071, country=China
[…]
[nklacd] organization=National Key Laboratory of Aircraft Configuration Design, city=Xi’an, postcode=710071, country=China
[…]
[seum] organization=School of Mathematics, Southeast University, city=Nanjing, postcode=211189, country=China
[…]
[itssky] organization=ITSSKY Technology Co., Ltd., city=Nanjing, postcode=210019, country=China
[…]
[ippt] organization=Systems Research Institute, Polish Academy of Sciences, city=Warsaw, postcode=01-447, country=Poland
[…]
[cityu] organization=Department of Data Science, City University of HongKong, city=Hong Kong, postcode=999077, country=China
[…]
[spu] organization=Suzhou Polytechnic University, city=Suzhou, postcode=215104, country=China
1 Introduction
Highways are the backbone of transportation networks in urban agglomerations, providing high-capacity mobility for the movement of people and goods between origins and destinations (Guan and He, 2008). The highway network is composed of basic road segments, entrance and exit ramps, and system interchanges (Manual, 2000). As key nodes connecting different highways, interchanges enable traffic flow to achieve continuous and rapid transitions in multiple directions through complex ramp systems. Moreover, interchange structures vary widely, including trumpet, turbine, cloverleaf, and their variants (Garber et al., 1999). However, during peak hours, interchange traffic flow faces severe challenges, as a large number of vehicles performing route changes converge at these nodes, resulting in traffic bottlenecks. These bottlenecks not only reduce interchange efficiency but also propagate along the mainline, ultimately leading to regional congestion (Ma et al., 2024; Mu et al., 2025). Consequently, such congestion seriously undermines both public travel satisfaction (Higgins et al., 2018) and logistics efficiency (Hu et al., 2018). Therefore, to alleviate congestion and maximize highway capacity, it is essential to monitor traffic operations at interchanges and to design advanced and effective management strategies. And accurate ramp traffic flow prediction provides the foundation for refined traffic control (Ma et al., 2020; Liu et al., 2016). Moreover, precise flow prediction helps traffic authorities capture macroscopic demand and analyze the microscopic composition of mainline flow, thereby clarifying the major sources and destinations of vehicles. Furthermore, it supports the design of targeted control strategies and enables dynamic monitoring and management of the transportation network (Wang et al., 2022; Zhang et al., 2022; Han et al., 2020). In addition, ramp traffic flow is indispensable for building complex traffic simulation models and achieving accurate traffic state estimation (Kan et al., 2020).
In practice, however, ramp flow data are frequently unavailable in real time. Not all ramps are equipped with sensors, and some lack communication channels (Nohekhan et al., 2021). Communication failures may also cause partial or complete data loss. For missing ramp detector data, imputation methods can be applied (Duan et al., 2016). Nevertheless, research on estimating ramp flows at locations lacking sensors remains limited. Conventional approaches primarily rely on origin–destination (OD) matrix–based link flow estimation. However, the acquisition of real-time OD matrices is extremely challenging (Marzano et al., 2009). Machine learning (ML) methods also face significant limitations, as they require extensive datasets and are highly dependent on data quality (Shi et al., 2021; Pereira et al., 2022). To overcome these limitations, recent studies have adopted transfer learning frameworks, utilizing mainline traffic flow to estimate ramp flows despite the lack of historical ramp data (Zhang et al., 2023; Ma et al., 2025).
In recent years, with the widespread deployment of Electronic Toll Collection (ETC) gantry systems, high-resolution spatiotemporal traffic data on highways have become increasingly accessible (Tsung et al., 2020; Zhou et al., 2025). Although ETC gantries are rarely installed on ramps due to high costs (Kan et al., 2020), ramp flows can be estimated by matching vehicle records at upstream and downstream gantries (Zhou et al., 2025). However, this approach encounters two primary challenges. First, legal requirements for data privacy and the technical complexity of system integration prevent authorities from processing vehicle records in real time. Consequently, ramps experience a “real-time blind spot”: while historical ramp flows can be obtained for model training, real-time ramp flows are unavailable as input during prediction. Second, data silos among different highway operators impede cross-regional vehicle tracking (Pospisil and Fric, 2016). Even when mainline data are available, equipment failures, adverse weather, or transmission errors often cause missing or anomalous values. Therefore, developing methods to accurately estimate and predict ramp flows from such incomplete mainline data—flows that are inherently difficult to observe—is both a critical technical requirement for fine-grained highway traffic management and a research problem of significant theoretical and practical importance.
Artificial Intelligence (AI) and ML provide promising solutions. From classical time-series models and Support Vector Machines (SVMs) to advanced deep learning models such as Convolutional Neural Networks (CNNs) (Jiang and Luo, 2022), Recurrent Neural Networks (RNNs) (Tedjopurnomo et al., 2020), and Transformers (Wen et al., 2023), these methods have been widely applied to traffic prediction. For ramp flow estimation, studies have used Random Forests (Kan et al., 2020) and deep learning models for hourly-level predictions (Zhang et al., 2023). Their common goal is to capture dynamic spatiotemporal dependencies in traffic data. Nevertheless, ML- and DL-based approaches face several limitations. Most rely on synchronous ramp data as input, which is infeasible under “real-time blind spot” conditions. Many also assume complete datasets, overlooking the prevalence of missing data in real-world scenarios. Transfer learning can mitigate some of these issues but remains limited for complex interchanges. As it does not fully exploit ETC-based historical ramp data. Consequently, existing methods inadequately capture the spatiotemporal coupling between mainline and ramp flows. This highlights the need for a universal model capable of handling diverse interchange structures.
To address these gaps, we propose the Spatio-Temporal Decoupled Autoencoder (STDAE) for ramp flow reconstruction. The framework is inspired by Masked Autoencoders (MAEs) (He et al., 2022), but it introduces a cross-modal reconstruction pretraining task to bridge the ramp “real-time blind spot.” It operates in two stages: (1) mapping learning, where the model reconstructs historical ramp flows from mainline flows, forcing it to learn spatiotemporal mappings; and (2) downstream prediction, where the learned representations are transferred for accurate ramp flow forecasting. The framework requires only real-time mainline data, even if incomplete, making it highly practical. Its design not only addresses missing input data but also improves adaptability and generalization across diverse interchanges.
The contributions of this study are as follows:
  • 1.
    Cross-modal reconstruction framework: Uses accessible mainline data to reconstruct ramp flows, addressing the ramp “real-time blind spot.”
  • 2.
    Robust mask-and-predict mechanism: Enhances prediction accuracy under incomplete data conditions.
  • 3.
    Comprehensive real-world validation: Evaluates performance on three real datasets and multiple sampling intervals, demonstrating improved accuracy and cross-scenario applicability.
The rest of the paper is organized as follows. Section 2 reviews related work. Section 3 introduces the framework and methodology. Section 4 describes data and preprocessing. Section 5 presents results and discussion. Section 6 concludes the study.
2 Related work
2.1 Short-term traffic prediction
Short-term traffic flow prediction is an essential component of Intelligent Transportation Systems (ITS) (Zhao et al., 2019). Accurate traffic flow forecasting provides critical decision support for traffic management authorities and helps effectively prevent and mitigate congestion (Liu et al., 2024b). Early studies mainly relied on classical statistical methods, such as autoregressive integrated moving average (ARIMA) models (Hamed et al., 1995; Zhang, 2003; Williams and Hoel, 2003) and Kalman filter models (Okutani and Stephanedes, 1984). These approaches are structurally simple, easy to implement, and can provide reliable baseline predictions under stable traffic conditions. However, traffic flow is inherently nonlinear and non-stationary, forming a complex time series. Traditional statistical models struggle to capture its intrinsic stochastic fluctuations and complex spatiotemporal dependencies, resulting in limited prediction accuracy under dynamic traffic conditions (Xu et al., 2022).
To overcome these limitations, researchers have increasingly adopted machine learning approaches (Tedjopurnomo et al., 2020). Shallow methods, such as SVM (Chen et al., 2015), K-Nearest Neighbors (KNN) (Zhang et al., 2009), and hybrid models (Li et al., 2023), can flexibly capture nonlinear relationships without requiring strict stationarity assumptions and exhibit strong adaptability to noise and local traffic flow trends. Consequently, these methods achieve higher prediction accuracy under highly fluctuating or complex traffic conditions. With the rapid development of research, deep learning has gradually become the mainstream approach (Tedjopurnomo et al., 2020; Yin et al., 2022; Jiang and Luo, 2022). In particular, RNNs and their variants, especially Long Short-Term Memory (LSTM) networks (Hu et al., 2025) and Gated Recurrent Units (GRU) (Dai et al., 2019), have demonstrated outstanding performance in capturing temporal dependencies. Meanwhile, CNNs have been employed to extract spatial topological features and combined with RNNs to form hybrid models (Jiang and Zhang, 2018; Méndez et al., 2023).
To further characterize the complex topological structure of traffic networks, researchers have proposed spatiotemporal prediction models based on Graph Neural Networks (STGNN) (Wu et al., 2020a; Yu et al., 2018; Zhao et al., 2019; Zheng et al., 2024; He et al., 2025), enhancing the model’s ability to capture spatiotemporal dependencies. In recent years, attention mechanisms and Transformers have also been introduced for traffic prediction (Liu et al., 2024a; Jin et al., 2025), enabling dynamic focus on critical time periods and key road segments, thereby improving the robustness and accuracy of predictions.
It is noteworthy that the majority of these studies have been applied to urban arterials or freeway mainlines where data collection is relatively comprehensive. The development and validation of these methods heavily depend on accessible mainline traffic flow data. In contrast, ramp traffic flows, which are often sparsely instrumented and difficult to observe directly, have received relatively limited attention in the literature, representing a significant gap in current research.
2.2 Ramp flow prediction
Highway ramp traffic flow prediction is a specialized branch within the field of dynamic traffic flow forecasting. The core task is to estimate the number of vehicles that will enter or exit the freeway mainline via ramps over a given future time interval (Zhang et al., 2023). As noted in Section 1, obtaining ramp flow data is challenging. Thus, existing research has primarily focused on methods for estimating flows at ramps without sensors. Traditional approaches mainly rely on OD matrices to allocate link flows. However, the real-time updating of OD matrices is extremely difficult, making these methods insufficient for dynamic traffic management (Marzano et al., 2009). Some studies combine traffic flow models with OD matrices. But these approaches are typically computationally intensive, exhibit limited accuracy, and require extensive parameter calibration (Chang and Li, 2002).
In recent years, data-driven machine learning and deep learning methods have become mainstream. Nevertheless, a common limitation of these models is their dependence on a large amount of auxiliary information to compensate for missing data. For instance, Nohekhan et al. (Nohekhan et al., 2021) proposed a model that requires static or semi-static inputs such as the annual average daily traffic (AADT), road classification, and free-flow speed. Similarly, Kan et al. (Kan et al., 2020) relied on ramp capacity and peak demand parameters derived from the Highway Capacity Manual (HCM) (Manual, 2000). This strong dependence on auxiliary information not only limits the deployability of models in data-constrained scenarios but also makes them vulnerable to sudden data loss caused by equipment failures or adverse weather conditions.
To alleviate the reliance on training data, recent studies have introduced transfer learning into ramp flow estimation. The core idea of transfer learning is to transfer knowledge learned from a data-rich source domain (e.g., mainline) to a data-scarce target domain (e.g., ramps) (Zhuang et al., 2020), providing a feasible solution for ramps without historical data. Although traffic states differ across highway segments, they exhibit inherent correlations, creating a natural scenario for the application of transfer learning (Zhuang et al., 2020). In ramp flow prediction, transfer learning is mainly applied in two ways: feature-representation transfer, which reduces the discrepancy in marginal distributions between the source and target domains, and parameter transfer, which mitigates differences in conditional distributions. Combining these strategies can significantly improve prediction accuracy and generalization capability in the target domain (Zhang et al., 2023).
However, existing transfer learning methods encounter new challenges when applied to structurally complex interchange scenarios. First, most prior studies focus on ramps with simple structures. Interchanges, however, have diverse layouts and complex ramp systems. This makes it difficult to directly apply existing models. Second, the data environment has evolved. With the deployment of ETC gantries and other technologies, historical ramp flow data can now be obtained through vehicle path matching. As a result, ramps are no longer entirely “data-free.” Historical time-series data are available offline, although real-time data remain inaccessible. This change presents new challenges for conventional methods. Traditional transfer learning methods are designed for fully data-sparse scenarios. They usually rely only on mainline data and ignore the valuable historical ramp data provided by ETC systems. This limits the ability of models to capture the complex spatiotemporal dependencies between mainline and ramp flows. Therefore, the key challenge is to use limited historical data effectively while handling incomplete mainline inputs. This is essential for developing accurate predictive models for ramps in complex interchanges.
2.3 Pretraining framework for traffic prediction
Inspired by successful practices in the field of natural language processing (NLP), pretraining has become a key paradigm for improving model performance (Liu et al., 2023b). The Transformer architecture, due to its strong ability to handle long sequences, forms the core of such pretrained models (Vaswani et al., 2017). In this context, researchers have extended the idea of pretraining to time series modeling. Through self-supervised learning, models extract effective feature representations from large-scale historical data. This enhances the performance of downstream prediction tasks (Shao et al., 2022a; Chi et al., 2025; Gao et al., 2024). Among these methods, MAE (He et al., 2022) has become a mainstream pretraining strategy. The core task of MAE is to mask part of the observed data and reconstruct it using the remaining visible data. This allows the model to learn deep relationships within the data. Such a property makes MAE naturally suitable for handling missing data, providing a new approach to improving the robustness of traffic prediction models.
Existing pretraining frameworks have explored various model structures and masking strategies. Some studies directly apply the MAE paradigm to spatiotemporal data. For example, TST (Zerveas et al., 2021) uses a masked Transformer architecture to learn representations of multivariate time series. Other models, including STEP (Shao et al., 2022a), DGCN-PTL (Chi et al., 2025), and STDMAE (Gao et al., 2024), further integrate the intermediate representations generated by MAE into downstream predictors to improve performance. Notably, STDMAE proposes a spatial-temporal decoupled masking mechanism. It performs masking and reconstruction separately along the temporal and spatial dimensions. This allows the model to better capture spatiotemporal heterogeneity. To model temporal dependencies more precisely, TS-MAE (Liu et al., 2025) constructs a masked autoencoder based on neural ordinary differential equations (Neural ODEs). This enables continuous modeling of evolution patterns. TS-MAE also uses a stepwise masking strategy to enhance learning robustness. In addition, some studies incorporate graph structures into pretraining. PesNet (Fang et al., 2024) uses graph convolution operators to recover masked data. JointSTNet (Cai et al., 2024) introduces a spatial graph capsule module and a temporal gating module. It also combines an adaptive masking strategy to improve the extraction and learning of complex spatiotemporal relationships.
Although these studies demonstrate the advantages of pretraining in capturing long-term and diverse spatiotemporal features, most pretraining tasks focus on self-reconstruction using historical data. That is, they reconstruct missing parts of the same source traffic flow. This “same-modality” reconstruction paradigm works well for general prediction or imputation tasks. However, it cannot directly solve the problem in interchange scenarios. In these cases, data from one modality (mainline traffic) must be used to predict another modality (ramp traffic). To address this, we propose a new cross-modal reconstruction pretraining task. The model uses mainline traffic flow data to reconstruct ramp traffic flows. In this framework, the model can learn more intrinsic patterns between mainline and ramp flows. Moreover, masking can be applied to the mainline inputs. This simulates real-world scenarios with partial missing data and enhances the robustness of the prediction model.
In summary, although ramp flow prediction makes great progress, current models still have some limitations. To compare the strengths and weaknesses of different approaches more clearly, this study reviews ramp flow prediction methods together with recent pretraining frameworks. Traditional approaches mainly use traffic flow models and dynamic OD estimation. They can predict ramp flows under certain conditions. However, they require complete and real-time data, and the computation cost is high. Machine learning methods reduce the impact of missing data by using additional static or semi-static features. But they rely heavily on auxiliary information, which limits scalability. Transfer learning methods transfer knowledge from mainline flows to ramps. This improves prediction accuracy, but they remain insufficient in complex interchange scenarios. Pretraining approaches take a different path. They learn representations from large-scale historical data through self-supervised learning. These methods show strong robustness and generalization. They also perform well under missing data conditions.
Table 1: Comparative analysis of different traffic prediction models across multiple evaluation metrics
Model1 Target2 Data Conditions3 Modeling Paradigm4 Model Evaluation5 Reference IDS DI DT Type TDM SDM VM RTC SGA OD RFP M-S, R-S M-C, R-C RT T $\checkmark$ $\times$ SB $\checkmark$ L Chang and Li (2002) RF/GBM RFP M-S M-C, R-M RT ML $\checkmark$ $\times$ RW $\checkmark$ M Kan et al. (2020) DDA+MT RFP M-S M-C, R-M RT TL $\checkmark$ $\times$ RW $\checkmark$ M Zhang et al. (2023) TST MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H Zerveas et al. (2021) STEP NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Shao et al. (2022a) DGCN-PTL NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Chi et al. (2025) STDMAE NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Gao et al. (2024) TS-MAE MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H Liu et al. (2025) PreSTNet NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Fang et al. (2024) JointSTNet NFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Cai et al. (2024)
Note: 1 Model: Name of the traffic flow prediction model evaluated.
2 Target: Application domain from which the dataset is derived, including Ramp Flow Prediction (RFP), Mainline Flow Prediction (MFP), or Network-Wide Flow (NFP).
3 Data Conditions: Assessment of the dataset used, including Input Data Source (IDS), Data Integrity (DI), and Data Timeliness (DT).
4 Modeling Paradigm: Type of model employed, including Traffic Flow Model (T), Machine Learning (ML), Transfer Learning (TL), or pretraining (Pre), along with Temporal Dependency Modeling (TDM) and Spatial Dependency Modeling (SDM) capabilities.
5 Model Evaluation: Validation method used (VM), including real-world (RW) or simulation/benchmark (SB) data, Real-Time Capability (RTC), and Scenario Generalization Ability (SGA).
Abbreviations: RFP: Ramp Flow Prediction, MFP: Mainline Flow Prediction, NFP: Network-Wide Flow Prediction,
M-S: Mainline Short time series, M-LS: Mainline Long and Short time series, R-S: Ramp Short time series,
M-C: Mainline Complete data, R-C: Ramp Complete data, R: Robust to missing data,
RT: Real-Time, Pre-N: pretraining on Non-real-time data, Test-RT: Test on Real-Time data,
T: Traffic Flow Model, ML: Machine Learning, TL: Transfer Learning, Pre: Pretraining Model.
Table 1 summarizes representative methods across several dimensions, including data conditions, modeling paradigms, dependency modeling, and generalization ability. The comparison shows that pretraining frameworks are more effective in handling missing data and capturing complex spatiotemporal dependencies. This provides a solid foundation for the method proposed in this study.
3 Methodology
This section describes the proposed model in detail. The overall framework of the model is illustrated in Figure 1. The STDAE model consists of two components: a pretraining model and a downstream prediction model. In the pretraining stage, to more efficiently capture the spatio-temporal dependencies of traffic data, we design an STDAE structure. This structure comprises two mutually independent modules: a Temporal AutoEncoder (TAE) and a Spatial AutoEncoder (SAE). Both modules share the same architecture but focus separately on reconstructing features in the temporal and spatial dimensions, respectively, to generate temporal and spatial representations of ramp traffic flow states. In the downstream prediction stage, these representations are processed through a multi-layer perceptron (MLP) module and fused with the hidden layer of the prediction model to obtain the final forecast of ramp traffic flow.
3.1 Problem Statement
In this study, the short-term prediction of ramp traffic flow at interchanges is defined as a temporal prediction problem. Consider a hub interchange with a double-line cross, where the mainline consists of $N=8$ directions, including four upstream and four downstream directions for the two intersecting lines. There are $M=12$ traffic movements (ramps), with each ramp connecting the upstream and downstream mainline flows from two corresponding directions.
$\mathbf{V}_{t-(T-1):t}\in\mathbb{R}^{T\times N\times F}$ denotes the historical mainline traffic flow feature time series for $T$ consecutive time steps, where $F$ represents the number of features per observation point. The objective is to predict the ramp traffic flow $\mathbf{Y}_{t+1:t+S}\in\mathbb{R}^{S\times M\times I}$ for the next $S$ time steps, given the ramp traffic graph $\mathcal{G}$:
$\hat{\mathbf{Y}}_{t+1},\ldots,\hat{\mathbf{Y}}_{t+S}]=F(\mathbf{V}_{t-(T-1):t},\mathcal{G}),$ (1)
where $\hat{\mathbf{Y}}_{j}\in\mathbb{R}^{M\times I}$ represents the predicted ramp traffic flow at the $j$th time step. In the dataset used in this study, $I=1$, corresponding to traffic volume measurements.
To model the dependencies among ramps, we construct a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node represents a ramp, and edges represent spatial or functional correlations between ramps. The graph has $M$ nodes corresponding to the $M$ ramps, and its adjacency matrix $\mathbf{A}\in\mathbb{R}^{M\times M}$ encodes the pairwise relationships among ramps:
[…]
In this study, we assume a fully connected graph by default, i.e., every ramp node is connected to all other ramp nodes, allowing the model to capture interactions among all ramps. This $M\times M$ fully connected graph structure is critical for accurate ramp traffic flow prediction.
3.2 The pretraining model
During the pretraining stage, we do not adopt a conventional autoencoding scheme that encodes the input into hidden representations and then reconstructs the input itself. Instead, the model reconstructs the ramp traffic flow sequences by leveraging the long historical time series of upstream and downstream mainline traffic flows. Spatiotemporal decoupling is the key strategy for modeling the spatiotemporal relationships between mainline and ramp traffic flows. Its core idea is to separate spatial dependencies from temporal dependencies and extract their features independently, thereby capturing complex spatiotemporal dynamics more accurately. In addition, an optional masking module is incorporated to address cases where part of the mainline traffic flow data is missing.
Figure 1: STDAE Pretraining and Prediction Framework
3.2.1 Input Feature Encoding
The input feature encoding is divided into two parts, namely feature fusion and feature encoding. In the feature fusion stage, we combine the mainline attributes—including traffic flow, speed, time of day, and day of week—represented as $\mathbf{V}_{t-(T_{\text{long}}-1):t}\in\mathbb{R}^{T_{\text{long}}\times N\times F}$. These features are reorganized into $\mathbf{X}\in\mathbb{R}^{T_{\text{long}}\times M\times C}$, where the feature representation of each ramp is constructed by aggregating the features from the corresponding upstream and downstream mainline connections. It is worth noting that this feature fusion step is required in both the pretraining and prediction stages, ensuring consistent representation of ramp-related inputs throughout the entire modeling process.
In the feature encoding stage, the input feature tensor $\mathbf{X}\in\mathbb{R}^{T_{\text{long}}\times M\times C}$ is selectively processed through a temporal mask or a spatial mask. If the mainline traffic data is complete, no masking is applied. When the task involves missing mainline data in certain directions or time intervals, the corresponding mask is used to simulate these missing values. To reduce computational cost for long sequences $T_{\mathrm{long}}$, the sequence is divided into $P$ non-overlapping segments, each of length $L=T_{\mathrm{long}}/P$, yielding $\mathbf{Z}\in\mathbb{R}^{M\times P\times LC}$. A fully connected layer then maps $\mathbf{Z}$ to an embedding $\mathbf{E}_{p}\in\mathbb{R}^{M\times P\times D}$. It is first passed through a layer normalization and then added to a 2D sinusoidal positional encoding $\mathbf{E}_{\mathrm{pos}}$ (Wang and Liu, 2021) to obtain the final input embedding $\mathbf{E}\in\mathbb{R}^{M\times P\times D}$.
3.2.2 STDAE
The embedding $\mathbf{E}$, derived from the mainline traffic features, is fed into two parallel modules that share the same architecture but serve different objectives.
Spatial Reconstruction Module (SAE). This module extracts the spatial heterogeneity features of ramp and mainline traffic flows through a stack of Transformer encoder layers. By leveraging the interaction between mainline and ramp flows, the module is able to capture the underlying spatial dependencies across different ramps.
Encoding. To enable the model to focus on the spatial dimension, the input embedding $E$ is first transposed along the spatial ($M$) and ramp ($P$) dimensions. The transposed tensor is then fed into the spatial encoder. After encoding and layer normalization, the result is transposed back to the original dimension order to form the spatial representation $H^{(S)}\in\mathbb{R}^{M\times P\times D}$. This process can be expressed as:
$H^{(S)}=\text{LayerNorm}\big(\text{EncoderS}(E^{T})^{T}\big)$ (3)
where $E^{T}\in\mathbb{R}^{M\times P\times D}$denotes the transpose of $E$ along its last two dimensions.
Decoding. The spatial representation $H^{(S)}$ is first passed through a linear mapping layer and then transposed again to match the input requirements of the spatial decoder. The decoder output is normalized and passed through the final output layer, followed by a final transpose to generate the spatially reconstructed sequence $\hat{Y}^{(S)}\in\mathbb{R}^{M\times P\times L}$:
$\hat{Y}^{(S)}=\Big[\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderS}(\text{Linear}(H^{(S)})^{T}))\big)\Big]^{T}$ (4)
Temporal Reconstruction Module (TAE). This module shares the same architecture as SAE but focuses on capturing long-range temporal dependencies.
Encoding. In the temporal domain, the input embedding $E\in\mathbb{R}^{M\times P\times D}$ is directly fed into the temporal encoder. The output is normalized to obtain the temporal representation $H^{(T)}\in\mathbb{R}^{M\times P\times D}$:
$H^{(T)}=\text{LayerNorm}(\text{EncoderT}(E))$ (5)
Decoding. The temporal representation $H^{(T)}$ is linearly mapped and directly input into the temporal decoder DecoderT without any transposition. The decoder output is normalized and passed through the output layer to obtain the temporally reconstructed sequence $\hat{Y}^{(T)}\in\mathbb{R}^{M\times P\times L}$:
$\hat{Y}^{(T)}=\text{OutputLayer}\big(\text{LayerNorm}(\text{DecoderT}(\text{Linear}(H^{(T)})))\big)$ (6)
The main difference between the two modules lies in the dimension processing strategy. The SAE module reorganizes the spatiotemporal dimensions via a transpose operation, allowing the Transformer to model dependencies across ramps and mainline. In contrast, the T-AE module preserves the original dimension order, focusing on capturing long-range temporal dependencies.
Both modules are optimized by computing reconstruction losses between $\hat{Y}^{(S)}$ and $\hat{Y}^{(T)}$ and the true ramp sequences. SAE captures spatial heterogeneity among ramps, while TAE models long-range temporal patterns. By reconstructing ramp traffic based on mainline flow features, the model implicitly learns the spatiotemporal dependencies between ramps and the mainline. The mainline acts as the traffic backbone, influencing ramp inflows and outflows and reflecting traffic propagation and interaction. Consequently, the reconstructed ramp sequences encode both spatial correlations and temporal dynamics, providing complementary spatiotemporal representations for downstream ramp traffic prediction tasks.
3.3 Downstream prediction model
In the downstream prediction model, it is possible to seamlessly integrate the pretrained model STDAE. This operation is accomplished by taking the spatial and temporal representations generated by STDAE and adding them to the hidden representation of the predictor.
Specifically, a long sequence of $T_{\text{long}}$ time steps is first fed into the pretrained spatial and temporal encoders to generate the corresponding spatial representation $H^{(S)}$ and temporal representation $H^{(T)}$. Subsequently, the commonly used short-term input sequence $X_{t-(T-1):t}$ is fed into the downstream prediction model to obtain the hidden representation $H^{(F)}\in\mathbb{R}^{T\times M\times D}$, where $D$ is the dimensionality of the hidden representation. During the input stage, the same masking strategy as in the pretraining stage is applied to ensure consistency.
To align with the hidden representation $H^{(F)}$, we extract the last $T^{\prime}$ patches from $H^{(S)}$ and $H^{(T)}$ and reshape them into tensors $H^{\prime(S)}\in\mathbb{R}^{T\times M\times D}$ and $H^{\prime(T)}\in\mathbb{R}^{T\times M\times D}$, respectively. These representations are then projected to the target dimension $D^{\prime}$ using two independent two-layer multilayer perceptrons.
The final augmented representation $H^{(\text{Aug})}\in\mathbb{R}^{T\times M\times D}$ is obtained by summing the projected spatial and temporal representations with the downstream hidden representation, as follows:
$H^{(\text{Aug})}=\text{MLP}(H^{\prime(S)})+\text{MLP}(H^{\prime(T)})+H^{(F)}.$ (7)
Up to this point, the augmented representation $H^{(\text{Aug})}$ incorporates the short-term representation learned by the downstream prediction model itself and the long temporal dependency features provided by STDAE, thus effectively compensating for the limited ability of traditional models to capture long-term dependencies and significantly improving the overall performance and generalization of downstream spatio-temporal prediction tasks.
In this study, the downstream forecasting model adopts GWNet (Wu et al., 2019). STDAE learns structured global representations by separately modeling spatial and temporal features, without requiring additional masking or reconstruction steps. By integrating the STDAE representations into the hidden states as auxiliary information, the model can better capture long-range dependencies, enabling more accurate and stable predictions in complex spatiotemporal tasks.
4 Dataset and processing
This section provides a detailed description of the entire process of data collection, cleaning, and feature engineering used in this study. The overall workflow is illustrated in Figure 2. The purpose of this process is to transform raw ETC transaction records into high-quality time series data of mainline and ramp traffic states that can be utilized by the model.
4.1 Define the scope
To ensure the generality and applicability of the proposed method and model, three representative and complex interchanges located in the southern region of Jiangsu Province, China, are selected as the study areas for data collection and analysis, as shown in Table 2. These three interchanges differ in network structure and traffic functionality, thereby providing a comprehensive testbed for evaluating the effectiveness of the proposed algorithm. In Figure 2, the QiLin interchange is taken as an example to illustrate the complete data processing workflow in detail.
Table 2: Details of Selected Interchanges
Interchange Name Location Connecting Expressways Interchange Type Primary Function Map View QiLin Nanjing G2503 Nanjing Ring Exp.
G42 Shanghai-Chengdu Exp.
Single-Loop Cloverleaf Connects Nanjing’s urban core, eastern regions, and the route to Shanghai; handles major urban-intercity traffic conversion. [Uncaptioned image]
DanYangXinQu Zhenjiang G42 Shanghai-Chengdu Exp.
S35 Fuyang-Liyang Exp.
Three-Loop Cloverleaf Manages traffic conversion between the east-west Shanghai-Nanjing corridor and north-south routes to central/northern Jiangsu. [Uncaptioned image]
XueBu Changzhou G4221 Shanghai-Wuhan Exp.
G4011 Yangzhou-Liyang Exp.
Four-Loop Cloverleaf Facilitates traffic connections between the Shanghai-Nanjing corridor and the southern Jiangsu / northern Zhejiang regions. [Uncaptioned image]
4.2 Data collection
To comprehensively capture the traffic dynamics of the interchange, eight ETC gantries located on the upstream and downstream mainlines closest to the interchange center are selected as data collection points. Two types of data are collected: (1) basic gantry information and (2) ETC transaction records. Basic gantry information provides static attributes of each collection point, including the operating company, road segment information, gantry ID, gantry name, latitude and longitude coordinates, highway kilometer marker, and direction of travel (upstream or downstream). This information provides the geospatial foundation for subsequent calculations of segment distance and traffic record matching. ETC transaction records constitute the core dynamic data of this study. They contain the passage information of every vehicle observed at the selected gantries. Key fields include the desensitized license plate number, passage time with second-level precision, and gantry ID. These data serve as the raw basis for computing traffic volumes, speeds, and ramp flows.
Figure 2: Data processing workflow, including the determination of the data collection scope, data collection, data cleaning, and feature extraction.
4.3 Data cleaning
The raw ETC transaction records are discrete event logs, which cannot be directly used for time series analysis. Therefore, the following data processing and feature engineering procedures are designed to transform the raw records into structured traffic features. First, the format of the acquired data is validated to ensure that all records follow a consistent structure and contain complete fields, which lays the foundation for automated processing. Erroneous records, such as timestamps outside the valid range or duplicate records of the same vehicle within the same second, are then removed. Next, each ETC record is matched with the corresponding gantry information using the gantry ID as the key index so that accurate geographical and road segment attributes can be assigned. Afterward, continuous event data are aggregated into discrete time series by defining the sampling granularity and splitting the study period into time windows. Mainline traffic volumes are computed by counting the number of vehicles passing through each gantry during every time window. To estimate mainline speed, the same license plate is matched across two adjacent gantries, the vehicle travel time over the segment is calculated, and the distance between gantries (obtained from kilometer markers) is divided by the travel time. The arithmetic mean of all vehicle speeds within a window is taken as the segment’s average speed. Finally, ramp flows, which are a key feature in this study, are inferred by matching license plates observed at upstream and downstream gantries, since ramps typically lack ETC gantries. Through this complete process, three core time series feature sets are obtained: mainline volume, mainline speed, and ramp volume.
In this study, the data covers the period from September 7, 2024, to September 29, 2024, totaling 23 days. Sampling intervals of 3 minutes, 5 minutes, and 10 minutes are adopted. A z-score normalization method is applied. The dataset is split into training, validation, and test sets in a ratio of 17:3:3, where the first 17 days are used as the training set, the middle 3 days as the validation set, and the last 3 days as the test set.
5 Experiments and results
This section presents a comprehensive evaluation of the proposed STDAE model on the datasets from Section 4. GWNet is selected as the downstream prediction model (Wu et al., 2019), forming the combined STAEGWNET model. Firstly, the overall performance of the STAEGWNET model is evaluated through comparison with several state-of-the-art models. Subsequently, ablation experiments are conducted to validate the effectiveness of the pretraining approach, followed by a deep analysis of the experimental results.
Table 3: QiLin—Overall Performance across Different Sampling Intervals
Model 3min 5min 10min Avg Rank $\downarrow$
MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$
GRU 5.66 $\pm$ 0.01 0.73 $\pm$ 0.01 8.79 $\pm$ 0.03 8.81 $\pm$ 0.04 0.71 $\pm$ 0.01 13.90 $\pm$ 0.07 17.20 $\pm$ 0.11 0.69 $\pm$ 0.01 26.89 $\pm$ 0.23 11.56
TCN 5.63 $\pm$ 0.02 0.73 $\pm$ 0.01 8.68 $\pm$ 0.03 8.82 $\pm$ 0.07 0.72 $\pm$ 0.01 13.91 $\pm$ 0.11 17.37 $\pm$ 0.24 0.71 $\pm$ 0.01 27.03 $\pm$ 0.51 12.11
AGCRN 5.02 $\pm$ 0.07 0.61 $\pm$ 0.01 7.92 $\pm$ 0.13 7.42 $\pm$ 0.10 0.55 $\pm$ 0.02 12.03 $\pm$ 0.18 14.04 $\pm$ 0.25 0.51 $\pm$ 0.01 22.73 $\pm$ 0.32 6.33
DCRNN 5.11 $\pm$ 0.04 0.64 $\pm$ 0.01 8.03 $\pm$ 0.08 7.55 $\pm$ 0.08 0.58 $\pm$ 0.01 12.36 $\pm$ 0.24 14.12 $\pm$ 0.24 0.55 $\pm$ 0.02 22.80 $\pm$ 0.52 9.44
MTGNN 6.53 $\pm$ 0.61 0.76 $\pm$ 0.09 9.88 $\pm$ 0.90 10.17 $\pm$ 0.98 0.71 $\pm$ 0.06 15.75 $\pm$ 1.87 19.12 $\pm$ 1.04 0.76 $\pm$ 0.07 28.78 $\pm$ 1.57 13.56
STPGNN 5.06 $\pm$ 0.01 0.67 $\pm$ 0.01 7.80 $\pm$ 0.03 7.32 $\pm$ 0.05 0.61 $\pm$ 0.01 11.59 $\pm$ 0.09 13.31 $\pm$ 0.05 0.56 $\pm$ 0.01 21.03 $\pm$ 0.12 5.89
TGCN 5.89 $\pm$ 0.09 0.72 $\pm$ 0.02 9.23 $\pm$ 0.18 8.97 $\pm$ 0.14 0.68 $\pm$ 0.02 14.25 $\pm$ 0.22 19.26 $\pm$ 0.26 0.75 $\pm$ 0.04 30.76 $\pm$ 0.50 12.78
STGCN 5.08 $\pm$ 0.02 0.61 $\pm$ 0.01 7.93 $\pm$ 0.04 7.45 $\pm$ 0.03 0.55 $\pm$ 0.01 12.09 $\pm$ 0.06 13.63 $\pm$ 0.11 0.51 $\pm$ 0.01 22.13 $\pm$ 0.23 6.67
D2STGNN 4.95 $\pm$ 0.04 0.61 $\pm$ 0.01 7.76 $\pm$ 0.09 7.34 $\pm$ 0.08 0.54 $\pm$ 0.01 12.02 $\pm$ 0.24 12.99 $\pm$ 0.12 0.48 $\pm$ 0.01 21.06 $\pm$ 0.24 3.56
GWNet 4.96 $\pm$ 0.03 0.62 $\pm$ 0.01 7.78 $\pm$ 0.06 7.18 $\pm$ 0.08 0.54 $\pm$ 0.01 11.72 $\pm$ 0.16 12.86 $\pm$ 0.21 0.48 $\pm$ 0.01 20.71 $\pm$ 0.40 2.89
STNorm 5.03 $\pm$ 0.03 0.62 $\pm$ 0.01 7.86 $\pm$ 0.08 7.33 $\pm$ 0.07 0.55 $\pm$ 0.01 11.79 $\pm$ 0.17 13.44 $\pm$ 0.25 0.50 $\pm$ 0.02 21.31 $\pm$ 0.41 4.89
STAEformer 5.10 $\pm$ 0.01 0.62 $\pm$ 0.01 7.98 $\pm$ 0.04 7.39 $\pm$ 0.09 0.55 $\pm$ 0.01 11.97 $\pm$ 0.20 13.24 $\pm$ 0.22 0.49 $\pm$ 0.02 21.59 $\pm$ 0.44 6.11
iTransformer 5.07 $\pm$ 0.01 0.62 $\pm$ 0.00 7.95 $\pm$ 0.01 7.72 $\pm$ 0.06 0.57 $\pm$ 0.01 12.31 $\pm$ 0.11 14.37 $\pm$ 0.15 0.54 $\pm$ 0.01 22.72 $\pm$ 0.33 8.22
STDAEGWNET 4.89** $\pm$ 0.02 0.60 $\pm$ 0.01 7.65** $\pm$ 0.04 7.01** $\pm$ 0.04 0.54 $\pm$ 0.01 11.29** $\pm$ 0.10 12.77 $\pm$ 0.13 0.47 $\pm$ 0.02 20.58 $\pm$ 0.25 1.00
Note: Use bold to indicate the best-performing model and underline to indicate the second-best model. * indicates $p<0.05$; ** indicates $p<0.01$ for significance test between the best and second-best models. Avg Rank shows the average ranking of each model across all sampling intervals and metrics.
Table 4: DanYangXinQu—Overall Performance across Different Sampling Intervals
Model 3min 5min 10min Avg Rank $\downarrow$
MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$
GRU 6.88 $\pm$ 0.02 0.63 $\pm$ 0.01 14.11 $\pm$ 0.07 10.61 $\pm$ 0.16 0.67 $\pm$ 0.01 23.00 $\pm$ 0.39 18.50 $\pm$ 0.30 0.77 $\pm$ 0.01 42.46 $\pm$ 0.74 11.67
TCN 6.96 $\pm$ 0.04 0.62 $\pm$ 0.01 14.26 $\pm$ 0.11 10.85 $\pm$ 0.16 0.66 $\pm$ 0.01 23.72 $\pm$ 0.44 18.91 $\pm$ 0.25 0.77 $\pm$ 0.01 43.15 $\pm$ 0.57 12.00
AGCRN 5.91 $\pm$ 0.12 0.51 $\pm$ 0.01 12.66 $\pm$ 0.23 8.90 $\pm$ 0.13 0.52 $\pm$ 0.02 20.04 $\pm$ 0.38 15.54 $\pm$ 0.44 0.56 $\pm$ 0.02 36.68 $\pm$ 1.10 5.56
DCRNN 6.13 $\pm$ 0.09 0.63 $\pm$ 0.02 13.09 $\pm$ 0.26 9.26 $\pm$ 0.14 0.61 $\pm$ 0.02 20.63 $\pm$ 0.40 16.40 $\pm$ 0.36 0.65 $\pm$ 0.03 38.38 $\pm$ 0.95 9.78
MTGNN 8.51 $\pm$ 0.53 0.80 $\pm$ 0.05 16.85 $\pm$ 0.89 12.97 $\pm$ 0.99 0.84 $\pm$ 0.09 28.40 $\pm$ 2.06 22.08 $\pm$ 1.31 0.85 $\pm$ 0.05 46.18 $\pm$ 2.04 14.00
STPGNN 5.79 $\pm$ 0.03 0.58 $\pm$ 0.01 12.33 $\pm$ 0.09 8.80 $\pm$ 0.08 0.58 $\pm$ 0.01 19.73 $\pm$ 0.23 14.69 $\pm$ 0.12 0.58 $\pm$ 0.01 34.58 $\pm$ 0.28 4.89
TGCN 6.93 $\pm$ 0.13 0.60 $\pm$ 0.02 14.58 $\pm$ 0.29 10.30 $\pm$ 0.25 0.66 $\pm$ 0.04 22.71 $\pm$ 0.41 19.26 $\pm$ 1.00 0.83 $\pm$ 0.08 45.61 $\pm$ 2.24 12.00
STGCN 5.89 $\pm$ 0.04 0.51 $\pm$ 0.01 12.56 $\pm$ 0.11 9.04 $\pm$ 0.10 0.51 $\pm$ 0.00 20.04 $\pm$ 0.27 15.61 $\pm$ 0.31 0.54 $\pm$ 0.01 36.45 $\pm$ 0.62 4.89
D2STGNN 5.77 $\pm$ 0.05 0.53 $\pm$ 0.01 12.40 $\pm$ 0.15 8.58 $\pm$ 0.14 0.50 $\pm$ 0.01 19.32 $\pm$ 0.32 14.65 $\pm$ 0.31 0.54 $\pm$ 0.03 34.87 $\pm$ 0.66 2.56
GWNet 5.66 $\pm$ 0.06 0.52 $\pm$ 0.02 12.17 $\pm$ 0.17 8.75 $\pm$ 0.09 0.55 $\pm$ 0.02 19.72 $\pm$ 0.30 15.01 $\pm$ 0.19 0.54 $\pm$ 0.01 35.54 $\pm$ 0.46 3.78
STNorm 5.93 $\pm$ 0.10 0.51 $\pm$ 0.01 12.70 $\pm$ 0.25 8.77 $\pm$ 0.12 0.53 $\pm$ 0.01 19.73 $\pm$ 0.35 15.30 $\pm$ 0.17 0.56 $\pm$ 0.01 36.00 $\pm$ 0.42 5.33
STAEformer 6.10 $\pm$ 0.05 0.54 $\pm$ 0.01 13.05 $\pm$ 0.15 9.05 $\pm$ 0.10 0.53 $\pm$ 0.02 20.31 $\pm$ 0.38 15.18 $\pm$ 0.54 0.60 $\pm$ 0.04 35.40 $\pm$ 1.18 6.89
iTransformer 6.14 $\pm$ 0.03 0.58 $\pm$ 0.01 13.24 $\pm$ 0.11 9.29 $\pm$ 0.10 0.60 $\pm$ 0.01 21.28 $\pm$ 0.30 16.03 $\pm$ 0.22 0.68 $\pm$ 0.02 38.33 $\pm$ 0.46 9.56
STDAEGWNET 5.61* $\pm$ 0.02 0.51 $\pm$ 0.01 12.07 $\pm$ 0.11 8.59 $\pm$ 0.08 0.53 $\pm$ 0.02 19.57 $\pm$ 0.25 14.69 $\pm$ 0.07 0.52* $\pm$ 0.01 35.37 $\pm$ 0.32 2.11
Note: Use bold to indicate the best-performing model and underline to indicate the second-best model. * indicates $p<0.05$; ** indicates $p<0.01$ for significance test between the best and second-best models. Avg Rank shows the average ranking of each model across all sampling intervals and metrics.
Table 5: XueBu—Overall Performance across Different Sampling Intervals
Model 3min 5min 10min Avg Rank $\downarrow$
MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ RMSE $\downarrow$
GRU 5.42 $\pm$ 0.03 0.64 $\pm$ 0.01 9.57 $\pm$ 0.09 8.30 $\pm$ 0.08 0.67 $\pm$ 0.01 15.13 $\pm$ 0.18 15.66 $\pm$ 0.18 0.87 $\pm$ 0.02 29.92 $\pm$ 0.48 12.78
TCN 5.37 $\pm$ 0.02 0.64 $\pm$ 0.01 9.44 $\pm$ 0.05 8.29 $\pm$ 0.07 0.67 $\pm$ 0.01 14.98 $\pm$ 0.13 15.15 $\pm$ 0.11 0.87 $\pm$ 0.02 28.63 $\pm$ 0.27 11.78
AGCRN 4.79 $\pm$ 0.07 0.58 $\pm$ 0.01 8.44 $\pm$ 0.11 7.37 $\pm$ 0.11 0.60 $\pm$ 0.02 13.42 $\pm$ 0.14 13.85 $\pm$ 0.51 0.74 $\pm$ 0.04 26.62 $\pm$ 1.07 8.67
DCRNN 4.86 $\pm$ 0.07 0.66 $\pm$ 0.02 8.38 $\pm$ 0.11 7.50 $\pm$ 0.19 0.64 $\pm$ 0.04 13.63 $\pm$ 0.53 13.73 $\pm$ 0.34 0.76 $\pm$ 0.04 25.47 $\pm$ 0.74 9.33
MTGNN 6.04 $\pm$ 0.36 0.67 $\pm$ 0.05 10.30 $\pm$ 0.90 8.70 $\pm$ 0.29 0.66 $\pm$ 0.04 14.74 $\pm$ 0.52 18.00 $\pm$ 0.94 0.94 $\pm$ 0.11 31.48 $\pm$ 2.96 13.56
STPGNN 4.74 $\pm$ 0.03 0.60 $\pm$ 0.01 8.26 $\pm$ 0.06 7.17 $\pm$ 0.05 0.59 $\pm$ 0.01 12.93 $\pm$ 0.10 12.86 $\pm$ 0.13 0.67 $\pm$ 0.02 24.21 $\pm$ 0.23 5.89
TGCN 5.38 $\pm$ 0.11 0.63 $\pm$ 0.02 9.65 $\pm$ 0.24 7.89 $\pm$ 0.24 0.65 $\pm$ 0.04 14.45 $\pm$ 0.57 14.99 $\pm$ 0.45 0.87 $\pm$ 0.06 28.87 $\pm$ 0.94 11.56
STGCN 4.73 $\pm$ 0.02 0.56 $\pm$ 0.01 8.33 $\pm$ 0.04 7.23 $\pm$ 0.07 0.55 $\pm$ 0.01 13.25 $\pm$ 0.11 13.13 $\pm$ 0.20 0.62 $\pm$ 0.03 25.11 $\pm$ 0.48 6.22
D2STGNN 4.62 $\pm$ 0.03 0.55 $\pm$ 0.01 8.17 $\pm$ 0.05 6.98 $\pm$ 0.07 0.50 $\pm$ 0.02 12.95 $\pm$ 0.27 12.38 $\pm$ 0.19 0.54 $\pm$ 0.04 24.15 $\pm$ 0.56 2.11
GWNet 4.64 $\pm$ 0.05 0.55 $\pm$ 0.02 8.23 $\pm$ 0.12 7.35 $\pm$ 0.14 0.54 $\pm$ 0.02 13.99 $\pm$ 0.47 12.34 $\pm$ 0.30 0.55 $\pm$ 0.02 24.18 $\pm$ 0.88 4.44
STNorm 4.69 $\pm$ 0.03 0.56 $\pm$ 0.01 8.31 $\pm$ 0.08 7.18 $\pm$ 0.20 0.53 $\pm$ 0.01 13.36 $\pm$ 0.53 12.73 $\pm$ 0.25 0.55 $\pm$ 0.02 25.09 $\pm$ 0.82 4.67
STAEformer 4.71 $\pm$ 0.04 0.56 $\pm$ 0.01 8.30 $\pm$ 0.07 7.12 $\pm$ 0.09 0.53 $\pm$ 0.02 13.01 $\pm$ 0.14 12.54 $\pm$ 0.14 0.60 $\pm$ 0.03 23.97 $\pm$ 0.29 4.22
iTransformer 5.10 $\pm$ 0.02 0.56 $\pm$ 0.01 9.10 $\pm$ 0.05 7.64 $\pm$ 0.05 0.54 $\pm$ 0.01 14.13 $\pm$ 0.11 13.68 $\pm$ 0.15 0.63 $\pm$ 0.02 26.35 $\pm$ 0.26 8.33
STDAEGWNET 4.58** $\pm$ 0.01 0.54* $\pm$ 0.01 8.11* $\pm$ 0.07 6.89** $\pm$ 0.04 0.52 $\pm$ 0.01 12.89 $\pm$ 0.14 12.30 $\pm$ 0.05 0.56 $\pm$ 0.02 23.70* $\pm$ 0.07 1.44
Note: Use bold to indicate the best-performing model and underline to indicate the second-best model. * indicates $p<0.05$; ** indicates $p<0.01$ for significance test between the best and second-best models. Avg Rank shows the average ranking of each model across all sampling intervals and metrics.
5.1 Experimental setup
Baselines. We compare STAEGWNET with four groups of representative benchmark models. The time-series models include GRU and TCN. The spatio-temporal GNN models comprise AGCRN (Bai et al., 2020), DCRNN (Li et al., 2017), MTGNN (Wu et al., 2020b), STPGNN (Kong et al., 2024), TGCN (Zhao et al., 2019), STGCN (Yu et al., 2018), and D2STGNN (Shao et al., 2022b). The spatio-temporal enhanced CNN models include GWNet (Wu et al., 2019) and STNorm (Deng et al., 2021). The spatio-temporal Transformer models consist of STAEformer (Liu et al., 2023a) and iTransformer (Liu et al., 2024a). These baselines span traditional statistical and machine learning methods to the latest graph neural networks and Transformer-based architectures, providing a comprehensive evaluation of the spatio-temporal modeling and forecasting capabilities of STAEGWNET. The input to these baseline models is consistent with STAEGWNET, consisting of mainline feature time series. In addition, we compare the results of HimNet (Dong et al., 2024), GWNet (Wu et al., 2019), and STDMAE (Gao et al., 2024) when using ramp flow time series as input with our model using mainline feature time series as input. The purpose is to verify whether using mainline features as input can achieve better prediction performance than directly using ramp flow sequences, thereby demonstrating the feasibility of predicting using only mainline data.
Settings. In the pretraining stage, the long temporal horizon is set to $T_{\text{long}}=1day$. In the forecasting stage, the input sequence length is fixed at $T=12$ time steps (Gao et al., 2024; Song et al., 2020) and the embedding dimension is $G=96$. The encoder comprises four Transformer layers and the decoder one Transformer layer, both employing four-head multi-head attention. To match the forecasting input, the patch size is $L=12$, and we set $T^{\prime}=1$, retaining only the last patch from the spatial representation $H^{(S)}$ and temporal representation $H^{(T)}$. Training uses the Adam optimizer with an initial learning rate of 0.002 and the mean absolute error (MAE) loss. Performance is evaluated by mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE). The results are averaged over 10 independent runs. All experiments are conducted on a Windows workstation equipped with a single NVIDIA V100 Tensor Core GPU. The experiments are conducted using the platform (Shao et al., 2024).
5.2 Overall performance
The performance of models is listed in Table 5, Table 5, and Table 5. The proposed STDAEGWNET model demonstrates significant superiority across all three interchange datasets. Moreover, for the 3-minute, 5-minute, and 10-minute forecasting tasks, the model achieves the best or second-best performance on the MAE, MAPE, and RMSE metrics, consistently maintaining the top overall average ranking.
Specifically, the model performance is reflected in the following four aspects. Overall Best Performance. The average ranking results further validate the comprehensive superiority of STDAEGWNET. The model average ranks on the QiLin, DanYangXinQu, and XueBu datasets are 1.00, 2.11, and 1.44, respectively, all ranking first. This performance clearly surpasses existing mainstream models, including graph neural network models, whose best ranks on the three datasets reach only 2.11, and Transformer-based models, whose best rank is only 4.22. This demonstrates that the proposed pretraining and downstream prediction framework can more effectively capture the spatiotemporal dependencies between mainline and ramp traffic, thereby exhibiting outstanding performance in practical prediction tasks. Significant Accuracy Advantage. In most prediction scenarios, STDAEGWNET achieves the lowest MAE, indicating that its predictions deviate least from the true values and outperform other baseline models in accuracy. At a 3 min sampling interval, the model mean absolute errors on the QiLin, DanYangXinQu, and XueBu datasets are 4.89, 5.61, and 4.58, respectively, all the lowest among the compared models and statistically significant. Even in scenarios where it does not achieve absolute optimality, performance remains close to the best, maintaining stable competitiveness. Strong Stability and Robustness. The root mean square error results show that STDAEGWNET generally outperforms other models, indicating concentrated prediction distributions and fewer extreme errors. On the QiLin dataset for 5 min predictions, STDAEGWNET achieves a root mean square error of 11.29, significantly lower than GRU at 13.90, TCN at 13.91, and DCRNN at 12.36. On the XueBu dataset for 10 min predictions, root mean square error is 23.70, lower than D2STGNN at 24.15 and GWNet at 24.18, demonstrating the model’s stability and robustness. Good Generalizability. Across three different hub datasets, STDAEGWNET consistently ranks among the top, indicating strong cross-scenario generalization and adaptability to diverse traffic characteristics. Compared with models that excel only on a single dataset, STDAEGWNET shows more robust overall performance. D2STGNN performs strongly on the XueBu and DanYangXinQu datasets with average ranks of 2.11 and 2.56 but performs relatively weaker on the QiLin dataset at 3.56. In contrast, STDAEGWNET maintains the best average ranks across all three datasets at 1.00, 2.11, and 1.44, highlighting superior robustness and consistency across scenarios.
Table 6: Performance Comparison with Advanced Ramp Flow Prediction Models Using Ramp Flow Sequences
Dataset Model 3min 5min 10min
MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE
QiLin HimNet 5.10 $\pm$ 0.04 0.63 $\pm$ 0.01 8.18 $\pm$ 0.09 7.74 $\pm$ 0.08 0.56 $\pm$ 0.01 13.08 $\pm$ 0.19 13.82 $\pm$ 0.33 0.45 $\pm$ 0.01 23.46 $\pm$ 0.79
GWNet 4.78 $\pm$ 0.04 0.59 $\pm$ 0.01 7.70 $\pm$ 0.09 7.20 $\pm$ 0.07 0.53 $\pm$ 0.01 11.87 $\pm$ 0.15 12.65 $\pm$ 0.20 0.45 $\pm$ 0.01 20.53 $\pm$ 0.31
STDMAE 4.66 $\pm$ 0.01 0.58 $\pm$ 0.00 7.36 $\pm$ 0.04 7.05 $\pm$ 0.03 0.53 $\pm$ 0.01 11.51 $\pm$ 0.07 13.23 $\pm$ 0.17 0.50 $\pm$ 0.01 21.28 $\pm$ 0.32
STDAEGWNET 4.891 $\pm$ 0.02 0.601 $\pm$ 0.01 7.651 $\pm$ 0.04 7.01a $\pm$ 0.04 0.541 $\pm$ 0.01 11.29a $\pm$ 0.10 12.771,3 $\pm$ 0.13 0.473 $\pm$ 0.02 20.581,3 $\pm$ 0.25
DanYang
XinQu
HimNet 6.10 $\pm$ 0.10 0.59 $\pm$ 0.02 12.97 $\pm$ 0.29 9.35 $\pm$ 0.18 0.56 $\pm$ 0.03 21.20 $\pm$ 0.55 15.72 $\pm$ 0.30 0.57 $\pm$ 0.04 37.85 $\pm$ 0.82
GWNet 5.75 $\pm$ 0.07 0.49 $\pm$ 0.01 12.35 $\pm$ 0.19 8.52 $\pm$ 0.12 0.49 $\pm$ 0.01 19.36 $\pm$ 0.38 14.51 $\pm$ 0.38 0.50 $\pm$ 0.02 35.40 $\pm$ 1.15
STDMAE 5.64 $\pm$ 0.02 0.48 $\pm$ 0.00 12.17 $\pm$ 0.05 8.52 $\pm$ 0.05 0.50 $\pm$ 0.01 19.30 $\pm$ 0.18 14.25 $\pm$ 0.13 0.50 $\pm$ 0.01 33.92 $\pm$ 0.35
STDAEGWNET 5.61a $\pm$ 0.02 0.511 $\pm$ 0.01 12.07a $\pm$ 0.11 8.591 $\pm$ 0.08 0.531 $\pm$ 0.02 19.571 $\pm$ 0.25 14.691 $\pm$ 0.07 0.521 $\pm$ 0.01 35.371 $\pm$ 0.32
XueBu HimNet 4.94 $\pm$ 0.01 0.57 $\pm$ 0.02 8.76 $\pm$ 0.03 7.26 $\pm$ 0.05 0.52 $\pm$ 0.01 13.34 $\pm$ 0.10 12.91 $\pm$ 0.16 0.55 $\pm$ 0.03 24.81 $\pm$ 0.26
GWNet 4.75 $\pm$ 0.02 0.54 $\pm$ 0.01 8.47 $\pm$ 0.05 7.27 $\pm$ 0.08 0.55 $\pm$ 0.01 13.43 $\pm$ 0.20 12.28 $\pm$ 0.15 0.53 $\pm$ 0.02 23.95 $\pm$ 0.46
STDMAE 4.68 $\pm$ 0.01 0.51 $\pm$ 0.00 8.38 $\pm$ 0.02 7.00 $\pm$ 0.01 0.49 $\pm$ 0.00 12.98 $\pm$ 0.05 12.28 $\pm$ 0.07 0.54 $\pm$ 0.01 23.50 $\pm$ 0.17
STDAEGWNET 4.58a $\pm$ 0.01 0.541 $\pm$ 0.01 8.11a $\pm$ 0.07 6.89a $\pm$ 0.04 0.522 $\pm$ 0.01 12.891,2 $\pm$ 0.14 12.301 $\pm$ 0.05 0.56 $\pm$ 0.02 23.701 $\pm$ 0.07
Note: Superscripts indicate significant improvements ($p<0.05$) of STDAEGWNET over: 1HimNet, 2GWNet, 3STDMAE, and aall three models.
Table 6 presents a comprehensive performance comparison between the proposed STDAEGWNET model, which uses mainline data, and three advanced benchmark models, HimNet, GWNet, and STDMAE, all of which rely on historical ramp data as input. The experimental results show that our proposed STDAEGWNET model, even without using the target ramp historical flow, achieves highly competitive performance and significantly outperforms models that directly use ramp historical data in many critical scenarios.
The results clearly demonstrate that STDAEGWNET consistently and significantly outperforms HimNet. Across nearly all evaluation metrics and scenarios, the improvements over HimNet are statistically significant ($p<0.05$), providing initial confirmation of the effectiveness of our model architecture. Furthermore, GWNet and STDMAE are two very strong benchmark models, making the comparison with them particularly valuable. The experiments reveal that in certain scenarios, models that directly leverage their own historical ramp data outperform STDAEGWNET. Nevertheless, STDAEGWNET still achieves superior performance in many cases. For example, in the 3 min forecasting tasks on the DanYangXinQu and XueBu datasets, its MAE and RMSE are significantly better than those of GWNet and STDMAE.
This comparative experiment highlights a central and important insight. Mainline traffic features contain richer and more predictive contextual information than the ramp’s own historical sequence. This seemingly counterintuitive result that indirect data can outperform direct data can be explained by the fact that ramp traffic flow largely reflects the outcome or manifestation of the broader traffic dynamics within the network. In contrast, mainline features such as flow and speed more effectively capture the underlying causes of ramp flow variations, such as upstream arrival waves or downstream capacity constraints. The STDAEGWNET model successfully learns these fundamental traffic evolution patterns from mainline data, which are indirect yet more informative, enabling it to deliver more accurate predictions than local perspective models that rely solely on the ramp’s own history.
In summary, the experimental findings strongly validate the effectiveness of the STDAEGWNET framework. It provides a practical solution to the challenge of unavailable ramp historical data under real-time blind spot conditions, and it achieves prediction accuracy that rivals or even surpasses advanced models relying on direct ramp history, thereby demonstrating dual advantages in both theory and practice.
5.3 Ablation study
Ablation Study on the Encoding Mechanism. To assess the effectiveness of the spatiotemporal encoding strategy, we design three model variants:
  • 1.
    TAE: Removes the spatial conditioning input and reconstructs sequences using only TAE, thereby gauging the contribution of spatial context.
  • 2.
    SAE: Removes the temporal conditioning input and reconstructs sequences using only SAE, thereby gauging the contribution of temporal context.
  • 3.
    w/o STDAE: Removes the temporal and spatial conditioning inputs.
Figure 3: Comparison of STDAE and its ablated versions on three datasets with 5-minute sampling granularity.
We report the experimental results on three datasets with a sampling granularity of 5 min. As shown in Figure 3, STDAE consistently outperforms all ablated versions. This advantage mainly stems from its design of reconstructing ramp traffic flow by leveraging mainline traffic features. Compared with approaches that directly rely on historical ramp flow, this mechanism extracts richer contextual information from mainline data, thereby compensating for the absence of ramp information under “real-time blind spot” conditions. In the spatial dimension, mainline flow reflects the traffic environment upstream and downstream of the ramp. In the temporal dimension, mainline features capture the underlying causes of ramp flow fluctuations, such as upstream arrival waves and downstream capacity constraints. Through this cross-source reconstruction mechanism, STDAE is able to learn the evolutionary patterns between ramps and mainlines, thereby significantly enhancing prediction accuracy and robustness.
Ablation Study on Predictor. To evaluate the generality of STDAE, we test five downstream predictors with different backbones, including RNN, GCN+RNN, GCN+TCN, and Transformer:
  • 1.
    STDAE-LSTM: Using LSTM as the predictor.
  • 2.
    STDAE-STGCN: Using STGCN as the predictor.
  • 3.
    STDAE-D2STGNN: Using D2STGNN as the predictor.
  • 4.
    STDAE-STAEformer: Using STAEformer as the predictor.
  • 5.
    STDAE-GWNET: Using GWNet as the predictor.
The experiments were conducted on three datasets with a 5 min sampling granularity. Table 7 presents the ablation results of STDAE across different downstream predictors. The experiments show that for all five downstream predictors, models enhanced with STDAE exhibit consistent and significant performance improvements across the three datasets.
Table 7: Ablation Study of STDAE Across Different Downstream Predictors
Model QiLin DanYangXinQu XueBu
MAE/MAPE/RMSE MAE/MAPE/RMSE MAE/MAPE/RMSE
LSTM $7.68/12.63/0.56$ $8.95/19.97/0.52$ $7.43/13.89/0.57$
STDAE-LSTM $\mathbf{7.50}^{*}/\mathbf{12.34}^{*}/\mathbf{0.55}$ $9.20/20.82/\mathbf{0.51}^{*}$ $\mathbf{7.22}^{*}/\mathbf{13.43}^{*}/\mathbf{0.54}^{*}$
STGCN $7.45/12.09/0.55$ $9.04/20.04/0.51$ $7.23/13.25/0.55$
STDAE-STGCN $\mathbf{7.37}^{*}/\mathbf{11.85}^{*}/0.56$ $\mathbf{8.94}^{*}/\mathbf{19.92}/0.53$ $\mathbf{7.09}^{*}/\mathbf{13.11}^{*}/\mathbf{0.54}$
D2STGNN $7.34/12.02/0.54$ $8.58/19.32/0.50$ $6.98/12.95/0.50$
STDAE-D2STGNN $\mathbf{7.21}^{*}/\mathbf{11.75}^{*}/\mathbf{0.53}$ $\mathbf{8.51}/\mathbf{19.26}/0.51$ $\mathbf{6.89}^{*}/\mathbf{12.77}/0.51$
STAEformer $7.39/11.97/0.55$ $9.05/20.31/0.53$ $7.12/13.01/0.53$
STDAE-STAEformer $\mathbf{7.26}^{*}/\mathbf{11.73}^{*}/0.55$ $\mathbf{8.80}^{*}/\mathbf{19.64}^{*}/0.54$ $\mathbf{7.11}/\mathbf{12.97}/0.54$
GWNET $7.18/11.72/0.54$ $8.75/19.72/0.55$ $7.35/13.99/0.54$
STDAE-GWNET $\mathbf{7.01}^{*}/\mathbf{11.29}^{*}/\mathbf{0.54}$ $\mathbf{8.59}^{*}/\mathbf{19.57}/\mathbf{0.53}$ $\mathbf{6.89}^{*}/\mathbf{12.89}^{*}/\mathbf{0.52}^{*}$
Note: Bold values indicate better performance (lower values) compared to baseline models. indicates statistically significant improvement ($p<0.05$).
Specifically, STDAE-LSTM shows reductions in MAE, MAPE, and RMSE compared to the baseline LSTM on the QiLin and XueBu datasets, indicating that the representations generated by STDAE effectively enhance the predictive capability of traditional sequential models. Similarly, when combined with graph neural network predictors, STDAE achieves superior performance across all datasets compared to their baseline versions, demonstrating that STDAE can extract rich spatiotemporal features from mainline traffic flows and convey them to downstream models. For more complex Transformer architectures, STDAE also brings noticeable performance gains. Overall, STDAE consistently and significantly improves performance across various predictors, validating that the mainline feature representations it generates are highly generalizable and robust.
These results indicate that STDAE not only independently enhances model performance but also serves as a general enhancement module, compatible with downstream predictors of different architectures, thereby improving the overall effectiveness of traffic flow forecasting.
5.4 Robustness on Missing Data
To evaluate the performance of the STDAEGWNET model under missing mainline data conditions, we introduced directional and temporal masking strategies to the mainline data to simulate data loss. The directional mask hides data from two mainline ETC gantries in the eastbound direction, while the temporal mask hides the last 6 steps out of every 12 time steps. Experiments were conducted on three datasets with sampling intervals of 3 min, 5 min, and 10 min. We compared models with and without STDAE, and the MAE results are shown in Figure 4. The results indicate that, across all datasets and sampling intervals, models incorporating STDAE significantly improve prediction accuracy. Compared to baseline models without STDAE, the STDAE-enhanced models consistently demonstrate clear advantages, achieving an average MAE reduction of approximately 2.23%. Moreover, the majority of the results passed the significance test ($p<0.01$).
Figure 4: MAE Comparison of STDAEGWNET with and without STDAE under Different Sampling Intervals and Significance Analysis. (a) 3 min. (b) 5 min. (c) 10 min. (d) Significance analysis.
Figure 5: Ramp-level Prediction of the QiLin Interchange with a 5 min Sampling Interval. (a) step = 3. (b) step = 6. (c) step = 12.
5.5 Accuracy Across Ramps
To further evaluate the model performance, we conducted a micro-level analysis of the prediction accuracy for individual ramps. This section aims to examine the improvement of STDAEGWNET over its backbone network, GWNet, in capturing the dynamics of specific ramp traffic flows. We selected the results of the QiLin interchange with a 5 min sampling interval and visualized the predictions for the first day in the test set.
As shown in Figure 5, the figure intuitively presents the comparison between the predicted curves and the ground truth for STDAEGWNET and GWNet at steps 3, 6, and 12. Visually, the prediction curves of STDAEGWNET closely follow the fluctuations of the actual traffic, particularly at the peaks and troughs of traffic flow, demonstrating stronger capability in capturing traffic dynamics.
For quantitative comparison, we further calculated the performance metrics for each turning ramp over the entire test set, and the detailed results are shown in Table 8. The quantitative results clearly confirm the conclusions from the visualization analysis. From the table, it can be seen that out of 12 ramps, STDAEGWNET outperforms the original GWNet in 10 ramps in terms of overall prediction performance. Particularly for major directions with higher traffic volumes and more challenging predictions, such as “W to E” and “E to W”, STDAEGWNET shows a significant advantage, with MAE reduced by approximately 9.0% and 4.9%, respectively, and RMSE also decreased accordingly. This indicates that the model predicts complex traffic patterns more accurately and robustly.
Table 8: Ramp-level Performance Comparison between STDAEGWNET and GWNet of the QiLin Interchange with a 5 min Sampling Interval
Direction STDAEGWNET GWNet Better Model
MAE MAPE (%) RMSE MAE MAPE (%) RMSE
E to N 2.90 53.29 3.84 2.84 50.42 3.78 GWNet
E to W 15.61 42.41 21.14 16.42 42.94 22.22 STDAEGWNET
E to S 5.64 35.75 8.12 5.82 34.84 8.38 STDAEGWNET
W to S 2.84 55.13 4.32 2.94 55.73 4.38 STDAEGWNET
W to E 14.11 66.34 19.55 15.50 70.78 21.99 STDAEGWNET
W to N 2.20 43.64 3.02 2.24 43.91 3.08 STDAEGWNET
S to E 9.79 63.65 13.08 9.87 63.00 13.54 STDAEGWNET
S to N 7.01 30.60 9.72 7.11 31.85 9.84 STDAEGWNET
S to W 4.34 88.80 6.30 4.58 86.21 6.69 STDAEGWNET
N to W 4.30 85.99 6.24 4.47 89.90 6.51 STDAEGWNET
N to S 6.13 28.84 8.50 6.39 29.05 8.98 STDAEGWNET
N to E 5.85 60.52 7.61 5.74 60.29 7.59 GWNet
It is worth noting that for the “E to N” and “N to E” ramps, GWNet exhibits a slight advantage. Analysis of the traffic characteristics reveals that these ramps likely have relatively low traffic volumes or simpler patterns, where the deep spatiotemporal context injected by STDAE pretraining does not play a critical role, and a strong direct prediction model can already achieve good performance. Overall, the ramp-level fine-grained comparative analysis strongly demonstrates the practical value of the STDAE pretraining framework. By learning deep mapping relationships between the mainline and ramps, it significantly enhances the backbone network performance at the micro-level, especially in handling complex and high-volume traffic scenarios, with substantial improvement in predictive accuracy.
6 Conclusion
This study proposes the STDAE framework, a two-stage approach that first performs cross-modal reconstruction pretraining and then integrates the learned representations with downstream predictors to tackle the problem of ramp flow prediction in detector-scarce highway interchanges. In the pretraining stage, STDAE reconstructs ramp flows from mainline traffic sequences collected by ETC systems, enabling its spatial and temporal branches to capture heterogeneous but complementary features. After feature fusion, the framework provides downstream predictors with rich long-range contextual information.
Comprehensive experiments on multi-week ETC datasets at three real-world cloverleaf interchanges and under multiple sampling intervals confirm the effectiveness of the proposed method. We adopt GWNet as the downstream predictor, and the combined model STDAEGWNET achieves the best overall performance across MAE, MAPE, and RMSE metrics among thirteen competitive baselines on all three datasets. Moreover, compared with advanced models that rely on historical ramp flow sequences, STDAEGWNET demonstrates comparable predictive capability. Since STDAE relies solely on widely available mainline ETC data, it enables fine-grained interchange control even in the absence of ramp detectors, without requiring hardware retrofitting. Extensive ablation studies further reveal that the proposed STDAE generates informative representations that effectively capture spatial and temporal dependencies between mainline and ramp flows. The learned embeddings are architecture-agnostic and can be seamlessly injected into other sequence or graph models with minimal modifications, serving as a plug-and-play enhancement for diverse forecasting pipelines. In addition, STDAE remains robust and achieves favorable performance even under scenarios with partially missing mainline data.
The current work focuses on multi-step short-term forecasting. Extending the framework to probabilistic prediction constitutes a natural next step. External factors such as incidents and weather are not explicitly modeled, and incorporating event-aware modules could further enhance robustness.
7 Acknowledgements
The authors gratefully acknowledge the support of ITSSKY Technology Co., Ltd. for providing the data used in this study. The authors would like to express their sincere gratitude to Dr. Wenbo Lu, Ph.D. candidate Feiyang Pan, Master’s student Tao Chen, Master’s student Qianyi Hu from Southeast University, and Ph.D. candidate Yibo Chen from East China University of Science and Technology for their invaluable support and thoughtful guidance. This research work is supported by the Big Data Computing Center of Southeast University.